An electric potential field is produced by point charges 1 mC and 4 mC located at (-2, 1, 5) and (1, 3, -1), respectively. The energy stored in the field is



1
Expert's answer
2021-07-07T12:52:50-0400

Gives

A(-2,1,5)and B (1,3-1)

"q_1=1mC,q_2=4mc"

"r_{AB}=\sqrt{3^2+2^2+(-6)^2}=7"

"U=\frac{kq_1q_2}{r_{AB}}"


"U=\frac{9\times10^9\times1\times10^{-3}\times4\times10^{-3}}{7}=\frac{36\times10^{3}}{7}J"

"U={5.14\times10^3}J"